OpenStudy (anonymous):

Determine the coordinates of the x-intercepts for the quadratic equation 5x2 – 21x + 4 = 0.

OpenStudy (anonymous):

5x² - 21x + 4 = 0 5x² - 20x - x + 4 = 0 5x(x-4) - (x-4) = 0 (5x-1) (x-4) = 0 x = 4, or x = 1/5